### v3

@@ 2 Related work
2.1 Active learning in computational scientific discovery
@@
Active learning techniques provide machine learning algorithms with more control over the training examples they use cohn_improving_1994 . A typical approach to active learning is via membership queries where the machine learner queries a data source to label instances according to some strategy. An active learner can be implemented to explore the version space which contains all hypotheses that are expressible by a hypothesis language and are consistent with a set of training examples mitchell_generalization_1982 . Instances are chosen according to a measure of informativeness such as entropy shannon_mathematical_1963 , diameter tosh_diameter_based_2017 , the size mitchell_generalization_1982 ; dasgupta_coarse_2005 and the shrinkage hocquette_how_2018 of the version space for eliminating competing hypotheses. For a binary classification task, when learners are allowed to look at $n$ unlabelled instances to compare, an active learner has $n$ times larger probability of selecting an instance with maximal entropy than a passive learner hocquette_how_2018 . The general principle for a binary active learner is to label instances closest to being consistent with half of the hypotheses in the version space mitchell_generalization_1982 . These instances reject up to half of the hypothesis space regardless of their classifications. In the optimal case when all instances are maximally discriminative, the target hypothesis can be found via a binary search with depth logarithmic of the hypothesis space size angluin_queries_2001 . $BMLP_{active}$ makes membership queries but our objective is to not only reduce the number of labels but also minimise the expected cost of experimentation.
@@
Computational systems such as the BACON systems langley_scientific_1987 , the LAGRANGE systems todorovski_integrating_2006 , recent equation discovery systems brunton_discovering_2016 ; guimera_bayesian_2020 ; petersen_deep_2020 are primarily designed to formulate symbolic hypotheses from experimental results. However, one cannot directly employ these systems for experimental planning. The automation of experimental design was investigated by the pioneering Robot Scientist King04:RobotScientist as an endeavour to bring together logical reasoning, active learning and laboratory automation. The Robot Scientist automatically proposes explanatory hypotheses, actively devises validation experiments, performs experiments using laboratory robotics, and interprets empirical observations. Abduction and active learning are supported by ASE-Progol ase_progol based on the ILP system Progol progol . Experiments were actively selected to minimise the expected cost of experimentation for learning gene functions in the aromatic amino acid synthesis pathway of yeast. It was combined with a laboratory robot to automatically reconstruct the yeast metabolic pathway model and the active learning strategy significantly outperformed random experiment selection King04:RobotScientist . Later, a high-throughput laboratory automation platform, Adam king_automation_2009 extended this approach by constructing a logical model of yeast consisting of 1200 genes. In comparison, we apply $BMLP_{active}$ to a genome-scale metabolic network, which is significantly larger than the models in King04:RobotScientist ; king_automation_2009 .
@@
Recently, large language models (LLMs) have been explored to automate experimental design via user-provided prompt texts boiko_autonomous_2023 ; bran_augmenting_2023 . LLMs are generative models of natural languages based on neural networks that are trained from textual data. In boiko_autonomous_2023 ; bran_augmenting_2023 , LLMs served as experiment planners and coding assistants. These platforms use LLMs for planning experimental steps using publicly available knowledge bases, writing and interpretation programs to control hardware in a laboratory environment and analysing collected experimental data. Their advantage is accessing rich information online and interacting with users while designing experiments. However, the consistency and faithfulness of their outputs cannot be guaranteed. It has been stressed that autonomous experimentation by AI should be employed with caution without the means to identify the source of error and explain observations ren_autonomous_2023 . While $BMLP_{active}$ cannot yet interact with biologists in natural language, interpreting and verifying its output are straightforward due to its logical representation and computation.
@@ 2.2 Learning genome-scale metabolic network models
GEMs contain known metabolic reactions in organisms, serving as functional knowledge bases for their metabolic processes. There has been an increasing interest in integrating machine learning methods with GEMs to improve the predictions of gene-phenotype correlations within biological systems angione_human_2019 ; sen_deep_2021 . This often requires constraint-based modelling techniques for simulating GEMs rana_recent_2020 . Flux balance analysis (FBA) palsson_systems_2015 is the most common mathematical approach to solve flux distributions in genome-scale models under the steady-state condition. FBA uses linear programming to model fluxes in a chemical reaction network represented by a matrix that contains the stoichiometric coefficients of metabolites in reactions. Ensemble learning frameworks such as wu_rapid_2016 ; oyetunde_machine_2019 ; heckmann_machine_2018 obtain support vector machines, decision trees and artificial neural networks to represent hidden constraints between genetic factors and metabolic fluxes. These systems cannot identify inconsistencies between simulation results and experimental data to self-initialise constraints to aid learning rana_recent_2020 . On the other hand, mechanistic information that reflects whole-cell dynamics has been incorporated to address these concerns, for instance in artificial neural networks to tune their parameters yang_white_box_2019 . The hybrid approach in faure_neural-mechanistic_2023 embeds FBA within artificial neural networks based on custom loss functions surrogating the FBA constraints. Recent research also explores autoencoders for learning the underlying relationship between gene expression data and metabolic fluxes via fluxes estimated by FBA sahu_advances_2021 . Currently, machine learning of genome-scale metabolic networks is significantly limited by the availability of training data sen_deep_2021 . In contrast, $BMLP_{active}$ does not use FBA so it is not dependent on the metabolic flux constraints. $BMLP_{active}$ also performs active learning to reduce the cost and the number of experiments. This is a notable advantage considering finite experiment resources and insufficient experimental data.
@@ 2.3 Matrix algebraic approaches to logic programming
In our BMLP framework, we use boolean matrices to evaluate fixpoints of large recursive datalog programs. Obtaining the least Herbrand model of recursive datalog programs can be reduced to computing the transitive closure of boolean matrices peirce_collected_1932 ; copilowish_matrix_1948 . Fischer and Meyer fischer_boolean_1971 study a divide-and-conquer boolean matrix computation technique by viewing relational databases as graphs. A similar approach is explored by Ioannidis ioannidis_computation_1986 for computing the fixpoint of recursive Horn clauses. An ILP system called DeepLog muggleton_hypothesizing_2023 employs boolean matrices for learning recursive datalog programs. It repeatedly computes square boolean matrices to quickly choose optimal background knowledge for deriving target hypotheses. The mapping from biological knowledge bases to propositional logic programs is explored in sato_boolean_2021 ; sato_differentiable_2023 where logic programs are represented as boolean matrices by considering gene regulatory networks as boolean networks. In contrast, we encode a GEM as a Petri net and a recursive datalog program for computing reachability queries.
@@
Recent work primarily studies the mapping of logic programs to linear equations over tensor spaces. SAT problems are investigated in lin_satisability_2013 under a linear algebra framework. Truth values of domain entities, logical relations and operators for first-order logic can be evaluated using tensors-based calculus grefenstette_towards_2013 . Computing algebraic solutions can approximate recursive relations in first-order Datalog sato_linear_2017 . Based on this approach, abduction can be performed by encoding a subset of recursive datalog programs in tensor space sato_abducing_2018 . Tensor representations allow differentiable infrastructures such as neural networks to perform probabilistic logical inferences cohen_tensorlog_2020 . This can be done using dynamic programming to propagate beliefs for stochastic first-order logic programs. The probabilistic parameters can be optimised using deep learning frameworks yang_differentiable_2017 . Limitations have been identified for solving certain recursive programs due to the difficulty of computing some arithmetic solutions whereas these can be solved by iterative bottom-up evaluations sato_linear_2017 . In addition, algebraic approximations are usually calculated in linear algebraic methods where the correctness of the solutions cannot be guaranteed. However, our BMLP approach does not have these issues since it uses boolean matrices for iterative bottom-up computation.
2.4 Petri nets in logic programming
Little attention has been paid to representing and simulating Petri nets with logic programs. With logic programming, techniques to execute, analyse and learn Petri nets can be implemented in a single programming language domenici_petri_1990 . The early work domenici_petri_1990 describes places and transitions of a Petri net in terms of n-ary ground facts. Srinivasan et al. srinivasan_knowledge_2012 attempt to learn n-ary ground facts that represent Petri net transitions in the ILP ILP1991 system Aleph aleph . Few studies have looked at encoding Petri nets by Answer Set Programming (ASP) gelfond_stable_1988 . Behrens and Dix behrens_model_2007 use Petri nets for analysing interactions in multi-agent systems with ASP programs. Anwar et al. anwar_encoding_2013 provide an ASP encoding of Petri nets with coloured tokens to model the behaviours of biological systems. This encoding is adapted by Dimopoulos et al. dimopoulos_encoding_2020 for simulating reversible Petri nets that allow transitions to fire in both directions. These systems all use high-level symbolic representations of Petri nets as deductive knowledge bases. Logic programming approaches rely on navigating background knowledge to determine the consistency of hypotheses. However, non-high-throughput deductive inferences limit these systems’ applicability for dealing with increasingly complex knowledge bases such as GEMs. $BMLP_{active}$ addresses this limitation by utilising a boolean matrix encoding of GEMs and boolean matrix operations to accelerate inferences.
3 Preliminaries
3.1 Matrix algebra for datalog
We introduce datalog programs and concepts relating to our BMLP framework based on LP_book ; ILP_foundation . A variable is a character string beginning with an uppercase letter. A function, predicate or constant symbol is a character string starting with a lowercase letter. A term is either a variable, a constant symbol, or a function symbol attached to a tuple of terms. A ground term contains no variables. An atom is a predicate symbol applied to a tuple of terms e.g. $r(a,Y)$ where $r$ is a predicate symbol, $a$ is a constant symbol and $Y$ is a variable. An atom is ground when it contains only ground terms. A Horn clause is a disjunctive clause with at most one positive literal. A definite clause is a Horn clause with exactly one positive literal and a definite program is a set of definite clauses. A query is a user-posted Horn clause without any positive literal. A general logic program engine such as a Prolog interpreter can evaluate a query by carrying out a refutation proof of its negation via SLD resolutions (Selective Linear Definite clause resolution). A unit ground clause (fact) is a definite clause with no body literal and no variables. Datalog programs are definite programs without function symbols. A first-order definite clause has the form, $A_{0}\leftarrow A_{1},...,A_{n}$ where $A_{i}=r_{i}(x_{1},...,x_{k})$, $r_{i}$ is a predicate symbol, $x_{j}$ is a term. The arity of a function or predicate symbol is its number of arguments, e.g. $r_{i}/k$ indicates $r_{i}$ has arity $k$. The literal $A_{0}$ is called the head of the clause and the literals $A_{i}$ are the body.
The Herbrand base $\mathcal{B_{P}}$ is the set of all unit ground clauses provable from the predicate and constant symbols in a first-order definite program $\mathcal{P}$. The Immediate Consequence Operator $\mathcal{T_{P}}$ van_emden_semantics_1976 is a mapping over subsets of $\mathcal{B_{P}}$ defined as:
$\forall I\subseteq\mathcal{B_{P}},\mathcal{T_{P}}(I)=\{\alpha\in\mathcal{B_{P}}\,|\,\alpha\leftarrow B_{1},...,B_{m}\,(m\geq 0)$
[…]
$\mathcal{T_{P}}$’s least fixpoint is $\mathcal{T_{P}}{\uparrow^{\omega}}$ = $\cup_{n\geq 0}\mathcal{T_{P}}{\uparrow^{n}}$ where $\mathcal{T_{P}}{\uparrow^{0}}=\emptyset$ and $\mathcal{T_{P}}{\uparrow^{n}}=\mathcal{T_{P}}(\mathcal{T_{P}}{\uparrow^{n-1}})$. A clause entailed by $\mathcal{P}$ is written as $\mathcal{P}\models\alpha$. The minimal Herbrand model $M(\mathcal{P})$ contains all unit ground clauses that are entailed by $\mathcal{P}$ and $M(\mathcal{P})$ = $\mathcal{T_{P}}{\uparrow^{\omega}}$.
A recursive program has a body literal that appears in the head of a clause. Specifically, we focus on linear and immediately recursive datalog as simple recursive datalog programs. A linear and immediately recursive datalog program has a single recursive clause with the recursive predicate appearing in the body only once ioannidis_computation_1986 . This subset of recursive datalog programs can represent the relational structure of reaction pathways. This is further elaborated in Section 4. An example of a linear and immediately recursive datalog program involving one non-recursive predicate is:
$r_{2}(X,Y)\leftarrow r_{1}(X,Y).$
$r_{2}(X,Y)\leftarrow r_{1}(X,Z),r_{2}(Z,Y).$
A recursive datalog program $\mathcal{P}$ can be written as matrices for bottom-up evaluation ceri_datalog_1989 . When $\mathcal{P}$ is linear and immediately recursive, it can be described by linear equations sato_linear_2017 :
$\textbf{R}_{2}^{0}=\textbf{0}$
$\textbf{R}_{2}^{k}=\textbf{R}_{1}+\textbf{R}_{1}\textbf{R}_{2}^{k-1}$
where unit ground clauses $r_{1}$ and a recursive clause $r_{2}$ can be represented as matrices R1, R2 such that $(\textbf{R}_{k})_{i,j}$ = 1 if $\mathcal{P}\models r_{k}(c_{i},c_{j})$ for constant symbols $c_{i},c_{j}$ and $(\textbf{R}_{k})_{i,j}$ = 0 otherwise. Addition “+” between two boolean matrices is defined as (A + B)i,j = Ai,j $\lor$ Bi,j. Multiplication “$\times$” is defined as $(\textbf{A}\times\textbf{B})_{i,j}$ = $\bigvee_{k=1}^{n}$ Ai,k $\land$ Bk,j and is abbreviated as AB.
3.2 Elementary nets
Figure 2: A one-bounded elementary net (OEN). Places are ellipses. Transitions are rectangles. Arcs are marked with weights. A black dot is a token. The initial marking is $m(c_{1})=m(c_{2})=1$. Firing transition $t_{1}$ gives a new marking $\{c_{3},c_{4}\}$.
@@ 3.2 Elementary nets
Petri nets have wide applications in Computer Science reisig_understanding_2013 . Elementary nets rozenberg_elementary_1998 are a class of Petri nets. We summarise key concepts and notions of Petri net and elementary nets from reisig_understanding_2013 ; rozenberg_elementary_1998 . A Petri net contains two disjoint sets of nodes: places $P$ and transitions $T$. The flow relation $F\subseteq(P\times T)\cup(T\times P)$ describes arcs between nodes. All places can be marked by tokens which symbolise the resources distributed in the network. The assignment of tokens to places is called a marking (of tokens). For a transition $t$, the pre-set ${}^{\bullet}t$ = $\{p\in P\,|\,pFt\}$ contains places whose tokens would be consumed by $t$ and its post-set $t^{\bullet}$ = $\{p\in P\,|\,tFp\}$ contains places that would receive tokens from $t$. Markings represent the states of a Petri net. The weight of an arc $(p,t)$ or $(t,p)$ is conventionally written as $\overline{pt}$ or $\overline{tp}$. Elementary nets are basic Petri nets where transitions consume one token from the pre-set and relocate a token to the post-set. In an elementary net, an arc weight $\overline{pt}=1\text{ if }(p,t)\in F$ and $\overline{tp}=1\text{ if }(t,p)\in F$. An elementary net is a tuple $(P,T,F)$.
@@
A one-bounded elementary net (OEN) is a class of elementary nets that has places marked with at most one token reisig_understanding_2013 . A marking $m$ in OEN is a characteristic function $m:P\to\{0,1\}$ defined as:
@@
4 Representing a GEM as a datalog program
4.1 Modelling a GEM with an OEN
Figure 3: OENs representing the qualitative relationship “$H_{2}$ and $O_{2}$ produce $H_{2}O$”. (a) Each place has up to one token. Each token represents the abundance of a chemical and this marking means $H_{2}$ and $O_{2}$ are available. (b) This model qualitatively describes that if $H_{2}$ and $O_{2}$ are available, then $H_{2}O$ becomes available by firing the transition.
We discuss the relationship between GEMs and OENs. Petri nets are commonly used formalism for representing complex biological systems such as genome-scale metabolic network models and diagnosing their dynamic behaviours sahu_advances_2021 . A genome-scale metabolic network model contains biochemical reactions of substances. Reactants are substances initially involved in a reaction and the products are the chemicals yielded by a reaction. A reaction is reversible when the conversion of reactants to products and products to reactants can take place simultaneously. Reversible reactions are commonly treated as two reactions in opposite directions. A reaction involves chemical substances $x_{i}$ and $y_{j}$:
[…]
A GEM is an OEN in the sense that a reaction is a transition, and reactants and products are places. Figure 3 presents a basic OEN modelling the water synthesis reaction involving $H_{2}$, $O_{2}$ and $H_{2}O$.
Given an auxotrophic mutant experiment, we model a GEM as an OEN to find the set of synthesisable substances by following the metabolic reaction pathways. This requires computing the necessary set of places for transitions to fire. Transitions need to be enumerated to examine if their necessary conditions can ever be met for them to fire. Every simulation of GEM involves a definition of the growth medium consisting of essential chemicals for cellular metabolism such as amino acids, inorganic salts and carbon sources. The definition of a growth medium serves as the basis of metabolic processes and gives an initial marking of OEN. Reachable places extending this marking are additional synthesisable substances.
4.2 Transforming an OEN into datalog
Each auxotrophic mutant experiment is treated as a datalog query to find all reachable places based on an initial marking. We evaluate queries of OEN reachability by transforming the OEN into a linear and immediately recursive datalog program. A datalog program consisting of clauses with arity two can be considered a directed graph where edges joining nodes represent the relation between terms. However, the multiplicity of the involved entities in a relation cannot be expressed this way. Instead, places and transitions are represented as constant and predicate symbols. Each argument of our transformed datalog program is a concatenation of constant symbols, e.g. $h2\_o2$. For our specific purpose, since all place nodes in the OEN are related to metabolites, every new concatenated constant symbol describes a set of metabolic substances.
This transformation with constant symbol concatenation generalises from a directed graph into a hypergraph berge_hypergraphs_1989 , a directed graph where an edge can connect to any number of nodes. Formally, a hypergraph $G=(V,E)$ is a tuple containing a set of nodes $V$ and a set of edges $E\subseteq 2^{V}\times 2^{V}$ between arbitrary subsets of $V$. Hypernodes are sets of nodes connected by such edges. The incidence graph of a hypergraph is bipartite in that for $v\in V$ and $e\in E$, $e$ is connected to $v$ if and only if $v$ and $e$ are nodes connected by an edge in the bipartite graph. Given an OEN, we consider a hypergraph with an incidence graph of the same bipartite structure. A hypernode represents a set of place nodes and is created for every transition’s input places from the OEN. Hypernodes are denoted by the new constant symbols that concatenate the member places’ names, e.g. a hypernode $\{H_{2},O_{2}\}$ is represented by the constant symbol $h2\_o2$. The number of hypernodes we create is fixed and limited by the number of transitions. This is because each reaction has a finite number of reactants and products. The dynamic behaviour of an OEN is specified via an initial marking. This mapping between an OEN and a hypergraph allows us to write a datalog program. A recursive clause is added which describes a path in the hypergraph.

### v4

@@ 7 Related work
7.1 Active learning
@@
A typical approach to active learning is via membership queries, where the machine learner queries a data source to label instances according to some strategy. An active learner can be implemented to explore the version space which contains all hypotheses that are expressible by a hypothesis language and are consistent with a set of training examples mitchell_generalization_1982 . Instances are chosen according to a measure of informativeness such as entropy shannon_mathematical_1963 , diameter tosh_diameter_based_2017 , the size mitchell_generalization_1982 ; dasgupta_coarse_2005 and the shrinkage hocquette_how_2018 of the version space for eliminating competing hypotheses. For a binary classification task, when learners are allowed to look at $n$ unlabelled instances to compare, an active learner has $n$ times larger probability of selecting an instance with maximal entropy than a passive learner hocquette_how_2018 . The general principle for a binary active learner is to label instances closest to being consistent with half of the hypotheses in the version space mitchell_generalization_1982 . These instances reject up to half of the hypothesis space regardless of their classifications. In the optimal case when all instances are maximally discriminative, the target hypothesis can be found via a binary search with depth logarithmic of the hypothesis space size angluin_queries_2001 . $BMLP_{active}$ makes membership queries to simultaneously reduce the number of labels and the expected cost of experimentation.
7.2 Computational scientific discovery
@@
Computational discovery systems langley_scientific_1987 ; todorovski_integrating_2006 ; brunton_discovering_2016 ; guimera_bayesian_2020 ; petersen_deep_2020 are mostly responsible for formulating symbolic hypotheses from experimental results, which are not directly applicable for experimental planning. The automation of experimental design was investigated by the pioneering Robot Scientist King04:RobotScientist to bring together logical reasoning, active learning and laboratory automation. The Robot Scientist automatically proposes explanatory hypotheses, actively devises validation experiments, performs experiments using laboratory robotics, and interprets empirical observations. It was combined with a laboratory robot to learn gene function annotations for aromatic amino acid pathways, and the active learning strategy significantly outperformed random experiment selection King04:RobotScientist . This was extended to identify genes encoding orphan enzymes from a logical model of yeast consisting of 1200 genes king_automation_2009 . In comparison, we apply $BMLP_{active}$ to a genome-scale metabolic network, which is significantly larger than the models previously examined. In addition, we examined isoenzymes, which were previously unexplored by logic programming due to larger hypothesis and experimental design spaces.
@@
Recently, large language models (LLMs) have been explored to automate experimental design via user-provided prompt texts boiko_autonomous_2023 ; bran_augmenting_2023 . LLMs are generative models of natural languages based on neural networks that are trained from textual data. In boiko_autonomous_2023 ; bran_augmenting_2023 , LLMs served as experiment planners and coding assistants. These platforms use LLMs to access publicly available knowledge bases, write and interpret programs to control laboratory hardware and analyse collected experimental data. The advantage is accessing rich information online and interacting with users while designing experiments. However, the faithfulness of their outputs raises concerns. It has been stressed that AI autonomous experimentation should be treated with caution, when there are no means to identify the source of error and explain observations ren_autonomous_2023 . $BMLP_{active}$’s logical representation and computation enhance faithfulness and output verifiability, which do not suffer from the ambiguity of natural languages.
@@ 7.3 Learning genome-scale metabolic network models
GEMs contain known metabolic reactions in organisms, serving as functional knowledge bases for their metabolic processes. There has been increasing interest in integrating machine learning methods with GEMs to improve the predictions of gene-phenotype correlations within biological systems angione_human_2019 ; sen_deep_2021 . This often requires constraint-based modelling techniques for simulating GEMs rana_recent_2020 . Flux balance analysis (FBA) palsson_systems_2015 is a widely used mathematical tool to solve flux distributions in genome-scale models under the steady-state condition. FBA uses linear programming to model fluxes in a chemical reaction network represented by a matrix that contains the stoichiometric coefficients of the metabolites in reactions. Ensemble learning frameworks such as wu_rapid_2016 ; oyetunde_machine_2019 ; heckmann_machine_2018 obtained support vector machines, decision trees and artificial neural networks to represent hidden constraints between genetic factors and metabolic fluxes. These systems cannot identify inconsistencies between the simulation results and the experimental data to self-initialise the constraints to aid learning rana_recent_2020 . On the other hand, mechanistic information that reflects whole-cell dynamics has been incorporated to address these concerns, for instance in artificial neural networks to tune their parameters yang_white_box_2019 . The hybrid approach in faure_neural-mechanistic_2023 embedded FBA within artificial neural networks based on custom loss functions surrogating the FBA constraints. Recent research also explored autoencoders to learn the underlying relationship between gene expression data and metabolic fluxes through fluxes estimated by FBA sahu_advances_2021 . Machine learning of genome-scale metabolic networks has been significantly limited by the availability of training data sen_deep_2021 . $BMLP_{active}$ does not use FBA, so it is not dependent on metabolic flux constraints. $BMLP_{active}$ also performs active learning to reduce the cost and number of experiments. This is a notable advantage considering finite experimental resources and insufficient experimental data.
@@ 7.4 Matrix algebraic approaches to logic programming
Obtaining the least Herbrand model of recursive datalog programs can be reduced to computing the transitive closure of Boolean matrices peirce_collected_1932 ; copilowish_matrix_1948 . Fischer and Meyer fischer_boolean_1971 studied a divide-and-conquer Boolean matrix computation technique by viewing relational databases as graphs. A similar approach was explored by Ioannidis ioannidis_computation_1986 for computing the fixpoint of recursive Horn clauses. An ILP system called DeepLog muggleton_hypothesizing_2023 employed Boolean matrices for learning recursive datalog programs. It repeatedly computes square Boolean matrices to quickly choose optimal background knowledge to derive target hypotheses. A propositional and Boolean matrices representation of biological knowledge bases was explored in sato_boolean_2021 ; sato_differentiable_2023 by considering gene regulatory networks as Boolean networks. In our BMLP framework, we use Boolean matrices to evaluate first-order recursive datalog programs.
@@
Recent work primarily studied the mapping of logic programs to linear equations over tensor spaces. SAT problems were investigated lin_satisability_2013 under a linear algebra framework, where truth values of domain entities, logical relations and operators for first-order logic are evaluated using tensors-based calculus grefenstette_towards_2013 . Computing algebraic solutions can approximate recursive relations in first-order Datalog sato_linear_2017 . Based on this approach, abduction can be performed by encoding a subset of recursive datalog programs in tensor space sato_abducing_2018 . Tensor representations allow differentiable infrastructures such as neural networks to perform probabilistic logical inferences cohen_tensorlog_2020 . This can be done using dynamic programming to propagate beliefs for stochastic first-order logic programs. The probabilistic parameters can be optimised using deep learning frameworks yang_differentiable_2017 . Limitations have been identified for solving certain recursive programs due to the difficulty of computing some arithmetic solutions, whereas these can be solved by iterative bottom-up evaluations sato_linear_2017 . In addition, algebraic approximations are usually calculated in linear algebraic methods where the correctness of the solutions cannot be guaranteed. However, our BMLP approach does not approximate Boolean matrices.
8 Conclusion and future work
To our knowledge, $BMLP_{active}$ is the first logic programming system able to operate on a state-of-the-art GEM, utilising the improved computational efficiency of Boolean matrices. $BMLP_{active}$ learned accurate gene function annotations while reducing the experimental costs by 90%. In addition, we demonstrated $BMLP_{active}$ actively learned gene-isoenzyme functions with 90% fewer training data points than random experiment selection. This suggests that $BMLP_{active}$ could offset the growth in experimental design space from genome-scale function learning. Future work could explore the use of Boolean matrix encoding of logic programs for even larger networks with GPU acceleration.
One limitation of $BMLP_{active}$ is that it cannot quantitatively predict the behaviours of biological systems. Future work could focus on using Probabilistic Logic Programming de_raedt_probabilistic_2015 and tensor-based algebra of logic programs sakama_logic_2021 given numeric training data. $BMLP_{active}$ also does not invent new predicates. Reusing invented predicates can reduce the size of a hypothesis learned by an ILP system and improve its learning performance ILP30 . Recent methods muggleton_hypothesizing_2023 ; evans_learning_2018 ; dai_abductive_2021 have enabled predicate invention in the context of matrix-driven ILP. While we touched on the theoretical properties of active learning, comparisons with other active learning approaches and theoretical claims can further contextualise our performance.
Importantly, we show that our approach can address the interactions between multiple genes, which are not addressable by current mutagenesis approaches. We aim to integrate $BMLP_{active}$ into a high-throughput experimental workflow wang_pooled_2018 for whole-genome function learning. Experimental data can first be compared with phenotypic predictions, narrowing the scope of the hypothesis space. With advanced biological gene editing tools, we can target specific gene edits to experimentally confirm candidate hypotheses during active learning.
Designing experiments to explain some hypothesised phenomenon requires deliberate thinking by scientists. A helpful scientific assistant could teach scientists how to create experiment plans with optimal outcomes. This aspect of human-AI collaboration has been studied in an emerging research area on Ultra-Strong Machine Learning (USML) US2018 ; Ai2021 ; sequential_teaching ; krenn_scientific_2022 . The present work can be extended to examine the effect of AI decision-making on human experiment selection and human comprehension of complex biological concepts.
Ethical Statement
The datasets used do not contain any personal information or other sensitive content. There are no ethical issues.
Data Availability
Source code and data are accessible on
https://github.com/lun-ai/BMLP_active/.
Acknowledgments
Appendix A One-bounded elementary net (OEN)
@@
Petri nets have wide applications in Computer Science [56]. Elementary nets [58] are a class of Petri nets. We summarise key concepts and notions of Petri net and elementary nets from [56, 58]. A Petri net contains two disjoint sets of nodes: places $P$ and transitions $T$. The flow relation $F\subseteq(P\times T)\cup(T\times P)$ describes arcs between nodes. All places can be marked by tokens which symbolise the resources distributed in the network. The assignment of tokens to places is called a marking (of tokens). For a transition $t$, the pre-set ${}^{\bullet}t$ = $\{p\in P\,|\,pFt\}$ contains places whose tokens would be consumed by $t$ and its post-set $t^{\bullet}$ = $\{p\in P\,|\,tFp\}$ contains places that would receive tokens from $t$. Markings represent the states of a Petri net. The weight of an arc $(p,t)$ or $(t,p)$ is conventionally written as $\overline{pt}$ or $\overline{tp}$. Elementary nets are basic Petri nets where transitions consume one token from the pre-set and relocate a token to the post-set. In an elementary net, an arc weight $\overline{pt}=1\text{ if }(p,t)\in F$ and $\overline{tp}=1\text{ if }(t,p)\in F$. An elementary net is a tuple $(P,T,F)$.
@@
A one-bounded elementary net (OEN) is a class of elementary nets that has places marked with at most one token [56]. A marking $m$ in OEN is a characteristic function $m:P\to\{0,1\}$ defined as:
@@
Appendix B Equivalence between OEN and datalog
Each auxotrophic mutant experiment is treated as a datalog query to find all reachable places based on an initial marking. We evaluate queries of OEN reachability by transforming the OEN into a linear and immediately recursive datalog program. A datalog program consisting of clauses with arity two can be considered a directed graph where edges joining nodes represent the relation between terms. Places and transitions are represented as constant and predicate symbols. Each argument of our transformed datalog program is a concatenation of constant symbols, e.g. $h2\_o2$. For our specific purpose, since all place nodes in the OEN are related to metabolites, every new concatenated constant symbol describes a set of metabolic substances.